Zegla manufacturing company has a payroll for the year in the amt of $4,588,655 with employee manual rate of $3.60/$100 of payroll, what is the premium in year 3 if the experience multiplier for year 2=1.62 and year 3=0.76

please show all work
Answer #1

Answer :

Payroll P=4588655

Manual rate M=3.60/100

Experience multiplier M1=1.62

Experience Multiplier M2=0.76

So premium =P*M*M1*M2=(4588655*3.6*1.62*0.76)/100=$203383.87